X- 32 Find the domain of the function: h(x) : 3. Give your answer in interval notation. x2-5x+6 X2-5x+b (x-3) (x+3)X
Answer #1

h(x) = 72 - 5x+6 > Va is real and x2 – 5x+6+ 0 =x20 and (x - 2)(x-3)=0 x20 and x + 2, X3 Domain: [0,2) (2,3) (3,0)